You sell short 200 shares of Doggie Treats Inc. which are currently selling at $52 per share. You post the 50% margin required on the short sale. If your broker requires a 28% maintenance margin, at what stock price will you get a margin call? (You earn no interest on the funds in your margin account and the firm does not pay any dividends)

You short-sell 217 shares of Alibaba, at $108 per share. If you wish to limit your maximum loss to $5,000, you should place a stop-buy order at ____. Ignore any dividends, trading costs, and margin interest.

VLC Inc. sold 200,000 shares in an initial public offering. The underwriter's explicit fees were $57 thousand. The offering price for the shares was $39, but immediately upon issue, the share price jumped to $55. What is the best estimate of the total cost to the firm of the equity issue?
Answer #1

1). Sale Proceeds from short sell = No. of shares * Current Price = 200 * $52 = $10,400

Amount deposited by investor = Sale Proceeds from short sell * Margin Requirement

= $10,400 * 0.50 = $5,200

Total Margin Requirement = Sale Proceeds + Amount deposited by investor

= $10,400 + $5,200 = $15,600

Let the stock Price = P

So, Cost of buying the asset = No. of shares * Shares Price = 200 * P

You will receive a margin call when the total required margin is higher than the total initial margin, i.e.,

Maintenance Margin = [Total Margin Requirement - Cost of buying the asset] / Cost of buying the asset

0.28 = [$15,600 - (200 * P)] / (200 * P)

0.28 * (200 * P) = $15,600 - (200 * P)

(56 * P) = $15,600 - (200 * P)

(56 * P) + (200 * P) = $15,600

(256 * P) = $15,600

P = $15,600 / 256 = $60.9375

2). Sale Proceeds from short sell = No. of shares * Current Price = 217 * $108 = $23,436

Let the stock Price = P

So, Cost of buying the asset = No. of shares * Shares Price = 217 * P

Loss = Cost of buying the asset - Sale Proceeds from short sell

$5,000 = (217 * P) - $23,436

217 * P = $5,000 + $23,436

217 * P = $28,436

P = $28,436 / 217 = $131.04

3). Underpricing Cost = [Current Share Price - Offer Price] * No. of shares

= [$55 - $39] * 200,000 = $16 * 200,000 = $3,200,000

Total Cost = Underpricing Cost + Underwriter's Fees

= $3,200,000 + $57,000 = $3,257,000